Sage Inc. experienced the following transactions for Year 1, its first year of operations:

  1. Issued common stock for $110,000 cash.
  2. Purchased $205,000 of merchandise on account.
  3. Sold merchandise that cost $162,000 for $322,000 on account.
  4. Collected $290,000 cash from accounts receivable.
  5. Paid $185,000 on accounts payable.
  6. Paid $64,000 of salaries expense for the year.
  7. Paid other operating expenses of $80,000.
  8. Sage adjusted the accounts using the following information from an accounts receivable aging schedule:
Number of Days Past Due Amount Percent Likely to Be Uncollectible Allowance Balance
Current $ 19,200 0.01
0–30 8,000 0.05
31–60 1,600 0.10
61–90 1,600 0.20
Over 90 days 1,600 0.50

Required

  1. Organize the transaction data in accounts under an accounting equation.
  2. Prepare the income statement, statement of changes in stockholders’ equity, balance sheet, and statement of cash flows for Sage Inc. for Year 1.
  3. What is the net realizable value of the accounts receivable at December 31, Year 1?
Answer #1
Event Assets = Liabilities + Equity Account
titles
Cash + AR - Allowance + Mer
Inv.
= Acc
Pay
+ Comm.
Stock
= Retained
Earnings
1. 110000 110000 NA
2. 205000 205000 NA
3a. -162000 -162000 Cost of goods sold
3b. 322000 322000 Revenue
4. 290000 -290000 NA
5. -185000 -185000 NA
6. -64000 -64000 Operating expenses
7. -80000 -80000 Operating expenses
8. 1872 -1872 Operating expenses
Bal. 71000 + 32000 - 1872 + 43000 = 20000 + 110000 + 14128 Net income

Sage Inc. Income Statement

$ $
Sales 322,000
Cost of Goods Sold 162,000
Gross Profit 160,000
Operating Expenses
Salaries Expense 64,000   
Other Operating Expenses    80,000
Bad Debt Expense 1872
Total Operating Expenses 145872
Net Operating Income 14128

Sage Inc. Statement of Changes in Stockholders' Equity

Beginning Common Stock 0
Issued during the period 110,000   
Ending Common Stock $ 110,000
Beginning Retained Earnings 0
Net income earned during the period    14128
Ending Retained Earnings 14128
Total Stockholders' Equity 124128

Sage Inc.Balance Sheet

Assets $ $
Cash 71,000
Accounts Receivable 32,000
Allowance for Doubtful Accounts (1872) 30128
Inventory 43,000
Total Assets 144128
Liabilities
Accounts Payable 20,000
Total Liabilities 20,000
Stockholders' Equity
Common Stock 110,000
Retained Earnings 14128
Total Stockholders' Equity 124128
Total Liabilities and Stockholders' Equity 144128   

Sage Inc. Statement of Cash Flows

$ $
Cash Flows from Operating Activities   
Cash collected from customers    290,000
Cash paid for inventories (185,000)   
Cash paid for operating expenses (144,000)
Net cash flows from Operating Activities (39,000)   
Cash flows from Investing Activities 0
Cash flows from Financing Activities
Issuance of Common Stock 110,000
Net cash flows from Financing Activities 110,000
Net change in cash 71,000
Beginning cash balance 0
Ending cash balance 71,000

net realizable value of the accounts receivable at December 31, Year 1 = 32000 - 1872 = 30128
Answer #2


image.png


image.png

image.png

image.png

image.png


image.png

image.png

answered by: CWF